OpenStudy (campbell_st):

If m> 0 find the value for m such that \[(424)_{m}= 3\times (123)_{m}\]

OpenStudy (blockcolder):

Assuming that 3 is in base 10, convert everything else to base 10: (note that m>4) \[4m^2+2m+4=3(m^2+2m+3)\\ 4m^2+2m+4=3m^2+6m+9\\ m^2-4m-5=0\\ (m-5)(m+1)=0\\ m_1=-1; m_2=5 \] Therefore, m=5.